### v1

@@ 1 Introduction
$\langle{\sf a}_{s+2\pi i}\,{\sf b}\rangle=\langle{\sf b}\,{\sf a}_{s}\rangle\,,$ (3)
@@
The KMS condition encodes a generalized notion of thermality Kubo:1957mj ; Martin:1959jp ; Haag:1996hvx . Indeed, when the modular flow aligns with temporal evolution, the KMS condition implies physical observers may experience thermal effects. Consider the Gibbs thermal state, where $\rho=e^{-\beta H}/Z$ and the modular flow corresponds to time-evolution in the Heisenberg picture:
@@
In this case, the KMS periodicity $s\sim s+2\pi i$ corresponds to periodicity in imaginary time, $t\sim t+i\beta$, the signature of thermality in QFT Matsubara ; Martin:1959jp .
@@
where $\xi(\mathbf{x})$ generates an infinitesimal local flow. In this case, $U(s)$ implements the flow along integral curves of $2\pi\xi(\mathbf{x})$. [^2] A more precise definition of geometric modular flows is the following Sorce:2020ama : Let $U$ be a unitary operator, and let $\varphi$ be a finite diffeomorphism. In particular, $\varphi=\varphi_{s}$ implements the flow with affine parameter $s$ along integral curves of $\xi$. Then, $U$ acts geometrically, implementing $\varphi$, if for any smearing function $f$ of any local operator $\phi(x)$, the operator $U^{{\dagger}}\phi[f]U$ can be written as $U^{{\dagger}}\phi[f]U=\phi[\tilde{f}]$, where $\mathrm{supp}(\tilde{f})=\varphi(\mathrm{supp}(f))$. The function $\tilde{f}$ need not equal the pullback $f(\varphi^{-1}(x))$; it can carry an extra conformal weight. In a CFT, if $\xi$ is a conformal Killing vector, the family of unitaries (2) built from eq. (5) satisfies these properties. Thus, we use eq. (5) as the defining property of a geometric modular flow. When these curves are future-directed, describing observer worldlines, the KMS condition (3) implies that observers may experience thermal effects. [^3] Note, we use $s$ to parameterize integral curves of $\xi$ rather than $2\pi\xi$. For this reason, we have replaced, the standard $i$-periodicity in the KMS condition with $2\pi i$ in eq. (3).
@@
Therefore, the number of states with geometric modular flow in any QFT is limited by the number of vector fields satisfying eq. (8), and depends on the conformal properties of the theory Sorce:2020ama . However, as noted in Jensen:2023yxy , it is also useful to consider a broader class of states whose modular flows are only instantaneously geometric, such that the modular Hamiltonian can be expressed by eq. (5) only on one Cauchy slice $\Sigma$ [^5] In CFT2, an instantaneously geometric flow can be extended to a geometric flow everywhere, since a vector field on an interval $A$ admits a unique extension to a conformal Killing vector field on the causal development of $A$. In this case, the condition (8) no longer applies.
It has recently been appreciated that gravitational constraints can be implemented to reduce the severity of ultra-local divergences in QFTs Witten:2021unn . In the language of von Neumann algebras, the type of the von Neumann algebra for a subregion is reduced from type III in continuum QFT to type II once the gravitational constraints are imposed Chandrasekaran:2022eqq ; Penington:2023dql ; AliAhmad:2023etg ; Klinger:2023tgi ; Kudler-Flam:2023qfl ; Jensen:2023yxy ; AliAhmad:2024eun ; Chen:2024rpx ; Kaplan:2024xyk ; DeVuyst:2024grw ; Kolchmeyer:2024fly ; Penington:2024sum . This type reduction implies the emergence of a well-defined trace that is unique up to rescaling, enabling precise definitions of density matrices and von Neumann entropies, which are unavailable in the type III setting. The outcomes of this research program have been striking, including a new semiclassical definition of the generalized entropy Witten:2021unn ; Jensen:2023yxy ; Chandrasekaran:2022cip . In this type-reduction procedure, it is important to consider states admitting geometric and instantaneously-geometric modular flows. Hence, our study of geometric modular flows and their abundance in CFTs is partially motivated by these results.
@@
$S(\rho_{\Psi}^{\mathcal{V}})=\alpha\int_{A}\text{d}\Sigma_{\mu}\,\frac{\xi^{\mu}_{\Psi}}{\norm{\xi_{\Psi}}^{d}}\,,$ (10)
where $\alpha$ is a state-independent thermal entropy density coefficient. This result is the covariant extension to higher dimensions of the $2d$ CFT expression in Wong:2013gua . We give various consistency checks of our formula in $d\geq 2$, including matching to several known examples. Notably, the physical reasoning which led us to propose eq. (10) is based on a local inverse temperature $\beta=2\pi\norm{\xi}$ associated instantaneously with observers along the flow, and thus we expect it applies also to states with only instantaneously geometric modular flows.
Secondly, we define a procedure in $d=2$ for removing the entanglement between the left and right Rindler wedges for some state $|\Psi\rangle$. When $|\Psi\rangle=|\Omega\rangle$, this procedure yields the Rindler vacuum $|\Omega\rangle_{\mathcal{R}}$ [^7] This is the vacuum state with respect to one-sided boosts, rather than time-translations. It is disentangled, such that we can write $|\Omega_{\mathcal{R}}\rangle=|0\rangle_{\mathrm{left}}\otimes|0\rangle_{\mathrm{right}}$. By contrast, the usual Minkowski vacuum is entangled: $|\Omega\rangle=\sum_{n}e^{-\pi E_{n}}|n\rangle_{\mathrm{left}}\otimes|n\rangle_{\mathrm{right}}/\sqrt{Z}$, where $E_{n}$ is the eigenvalue of $|n\rangle_{\mathrm{right}}$ under right-sided boosts.. We then propose the existence of analogous states in higher $d$, characterized by the following one-point function:
@@
Again, the physical reasoning which leads to this formula is based on local temperatures associated instantaneously with observers along the flow.
@@
Various technical details are provided in the appendices. In appendix A, we demonstrate that if two states with geometric modular flows are related by a conformal unitary, the flows simply transform as the push-forward under the corresponding conformal map, which is essential for our construction of states with geometric modular flows. In appendix B, we show that states with geometric modular flows in a subregion can be perturbed outside this subregion, without changing their flow inside, which is important for understanding the uniqueness of our various constructions. In appendix C, we provide supplementary calculations to ensure that the conformal mappings discussed in sections 2.1 and 2 are well-defined. In appendix D, we use the entanglement entropy formula discussed in Cardy:2016fqc to independently check the formula (61) in Wong:2013gua , which is a special case of our proposal eq. (10). Appendix E deals with nonlocal transformation rules for primary operators and the subtleties arising in the study of special conformal transformations in Minkowski spacetime. Finally, Appendix F shows that in the case of multiple intervals in CFT2, our proposed entropy formula (10) reproduces the correct entanglement entropy even when the modular Hamiltonian contains nonlocal terms. Readers seeking an introductory reference to modular flows and the theory of algebraic QFT are encouraged to consult Sorce:2023fdx ; Sorce:2024pte ; Sorce:2023gio ; Witten:2018zxz ; Hollands:2018wxj ; Hollands:2022dem ; Fewster:2019ixc ; Hollands:2014eia .
@@ 2 Excited states in $d=2$ with geometric modular flows
  • are future-directed in $\mathcal{V}$ and past-directed in $D(\bar{A})$, [^9] When Haag duality holds, we have $U_{\Psi}^{D(A)}(s)=U_{\Psi}^{D(\bar{A})}(-s)$ Haag:1996hvx . Thus, if the modular flow with respect to $D(\bar{A})$ is future-directed in $D(\bar{A})$, then the modular flow with respect to $\mathcal{V}$ is past-directed in $D(\bar{A})$.
@@ 2.1 Excited states in the vacuum sector
The conformal Killing equation (8) in $d=2$ has general solution $\xi=c(u)\partial_{u}+b(v)\partial_{v}$. We accordingly define a vacuum-sector Unruh flow with respect to $\mathcal{V}=\mathcal{R}$ as a flow of this form, which approaches the vacuum modular flow Bisognano:1976za
@@
Let such a conformal transformation $\varphi$ be implemented at the quantum level by a conformal unitary $U_{\varphi}$. For a precise discussion of such unitaries, see e.g. Fewster:2004nj . Under $U_{\varphi}$, conformal primary operators transform according to the first derivatives of $\tilde{u}(u)$ and $\tilde{v}(v)$:
@@
Moreover, the Schwarzian appearing in the transformation of the stress tensor contains a third derivative, so will require that $\tilde{u}(u)$ and $\tilde{v}(v)$ are at least three times differentiable, i.e., $\tilde{u}(u),\tilde{v}(v)\in C^{3}(\mathbb{R})$. This guarantees that the state $|\Psi\rangle:=U_{\varphi}|\Omega\rangle$ has well-defined expectation values for most primary and quasi-primary operators of interest. [^13] Since higher derivatives will appear in transformation laws for descendants, infinite differentiability would be ideal, but it is difficult to explicitly check it in the cases of interest. Some of the conformal maps considered in this paper are not smooth, notably the disentangling maps of section 3.4. See also section 2.
@@
$U_{\Psi}(s):=U_{\varphi}\,U_{\Omega}(s)\,U_{\varphi}^{\dagger}\,,$ (16)
because $U_{\Psi}$ yields the relevant KMS condition. Note, the KMS condition is satisfied also when the subregion is not preserved; see appendix A. The main point is that the modular flow for $|\Psi\rangle$ in $\mathcal{R}$ is a composition of geometric unitaries, and hence is geometric. [^14] Had we considered a highest-weight state $|h\rangle$ instead of the vacuum, the modular flow unitary $U_{\Psi}(s):=U_{\varphi}\,U_{h}(s)\,U_{\varphi}^{\dagger}$ (17) for $|\Psi\rangle:=U_{\varphi}|h\rangle$ has a geometric piece $U_{\varphi}$ and a generically non-geometric piece $U_{h}(s)$ (see, e.g., Caputa:2023vyr ).
@@
maps spatial infinity to the point $(\tilde{t}_{\cal{D}},\tilde{x}_{\cal{D}})=(0,R)$, such that $\mathcal{R}$ is mapped to $\mathcal{D}:=D(A)$, where $A$ is the interval $\tilde{x}_{\cal{D}}\in(-R,R)$ at $\tilde{t}_{\cal{D}}=0$. Since this map preserves the vacuum sector (see footnote 16), the pushforward of a vacuum-sector modular flow $\tilde{c}\,\partial_{\tilde{u}}+\tilde{b}\,\partial_{\tilde{v}}$ under this map is a vacuum-sector modular flow associated with the diamond $\mathcal{D}$. Since a generic state in the CFT is expected to have boost-like modular flow at $\partial A$, we expect
@@ 2.2 Excited states in the thermal sector
which exactly cancels eq. (37). This implies that there is the same amount of entropy in the two copies of the system, and thus the state is pure, as one can derive from eq. (69). As we derive below, in the thermal sector, the purity condition eq. (19) generalizes to
@@
These falloffs are more stringent than their analogue in eq. (13). [^18] Note that the arguments used in eq. (33) to fix the asymptotic behavior do not apply here, due to the absence of special conformal symmetry (see footnote 16). We excluded ${\cal O}(1)$ contributions, since the leading $\frac{\beta}{2\pi}$ term ensures the state reduces to the thermal state with inverse temperature $\beta$ at spatial infinity. We moreover excluded ${\cal O}(\frac{1}{\tilde{u}})$ contributions because otherwise, the resulting states would have infinitely more entanglement entropy in the Rindler wedge compared to $|\beta\rangle$, see section 3.2 for details. Note, however, that a similar analysis may apply for more general falloffs, as we discuss further in sections 3.2 and 6.
@@
What happens when an Unruh flow $\xi$ does not satisfy our purity condition? In that case, the vacuum modular flow can be transformed to $\xi$ using a conformal map $\varphi$ which fails to be differentiable at the origin. [^20] One can instead use a conformal map $\varphi$ which is not differentiable at infinity. Similar statements hold for flows not satisfying the boost-like property, such that $\varphi$ is non-invertible at the origin. See section 3.3. Due to its Schwarzian transformation rule (eq. (48) below), the resulting stress tensor is divergent along $\tilde{u}=0$ and $\tilde{v}=0$. These divergences propagate to other fields in the theory. [^21] Similarly, the Rindler vacuum stress tensor acquires divergences on the horizon due to the breaking of entanglement between left and right wedges Hartman:2015 ; Chen:2022 . The singular behavior resulting from violations of our purity condition reflects the idea that a pure state cannot be mapped to a mixed state by a unitary operator. [^22] For the vacuum sector, eq. (19) requires the entanglement entropy to be the same individually for both the left-moving and right-moving contributions. This indicates that our conformal unitaries cannot be used to introduce entanglement between right- and left-movers. Note that this is different from a shockwave state produced by the insertion of a primary operator $O_{1}$ at the origin; in this case, only primary operators with the same scaling dimension receive divergences at null separation.

### v2

@@ 1 Introduction
$\langle{\sf a}_{s}\,{\sf b}\rangle=\langle{\sf b}\,{\sf a}_{s+i}\rangle\,,$ (3)
@@
The KMS condition encodes a generalized notion of thermality Kubo:1957mj ; Martin:1959jp ; Haag:1996hvx . Indeed, when the modular flow aligns with temporal evolution, the KMS condition implies physical observers may experience thermal effects. Consider the Gibbs thermal state, $\rho=e^{-\beta H}/Z$, where the modular flow corresponds to time-evolution in the Heisenberg picture:
@@
In this case, the KMS periodicity $s\sim s+i$ corresponds to periodicity in imaginary time, $t\sim t+i\beta$, the signature of thermality in QFT Matsubara ; Martin:1959jp .
@@
where $\xi(\mathbf{x})$ generates an infinitesimal local flow. In this case, $U(s)$ implements the flow along integral curves of $2\pi\xi(\mathbf{x})$. [^2] A more precise definition of geometric modular flows is the following Sorce:2020ama : Let $U$ be a unitary operator, and let $\varphi$ be a finite diffeomorphism. In particular, $\varphi=\varphi_{s}$ implements the flow with affine parameter $s$ along integral curves of $\xi$. Then, $U$ acts geometrically, implementing $\varphi$, if for any smearing function $f$ of any local operator $\phi(x)$, the operator $U^{{\dagger}}\phi[f]U$ can be written as $U^{{\dagger}}\phi[f]U=\phi[\tilde{f}]$, where $\mathrm{supp}(\tilde{f})=\varphi(\mathrm{supp}(f))$. The function $\tilde{f}$ need not equal the pullback $f(\varphi^{-1}(x))$; it can carry an extra conformal weight. In a CFT, if $\xi$ is a conformal Killing vector, the family of unitaries (2) built from eq. (5) satisfies these properties. Thus, we use eq. (5) as the defining property of a geometric modular flow. When these curves are future-directed, describing observer worldlines, the KMS condition (3) implies that observers may experience thermal effects. [^3] Note, we use $s$ to parameterize integral curves of $\xi$ rather than $2\pi\xi$. For this reason, we must replace the standard $i$-periodicity in the KMS condition (3) with $2\pi i$ going forward. The Canadian authors apologize for this unconventional choice. In most equations, standard conventions can be restored via $\xi\mapsto\xi/2\pi$.
@@
Therefore, the number of states with geometric modular flow in any QFT is limited by the number of vector fields satisfying eq. (8), and depends on the conformal properties of the theory Sorce:2020ama . However, as noted in Jensen:2023yxy , it is often useful to consider a broader class of states whose modular flows are only instantaneously geometric, such that the modular Hamiltonian can be expressed by eq. (5) only on one Cauchy slice $\Sigma$ [^5] In the context of CFT2, it suffices to study globally geometric flows, because in $d=2$ any vector field on an interval $A$ admits a unique extension to a conformal Killing vector on the causal development of $A$. In this case, the condition (8) no longer applies.
Our study of geometric modular flows and their abundance in CFTs is partially motivated by certain results in semiclassical gravity. It has recently been appreciated that gravitational constraints can be implemented to reduce the severity of ultra-local divergences in QFTs Witten:2021unn . In the language of von Neumann algebras, the type of the von Neumann algebra for a subregion is reduced from type III in continuum QFT to type II once the gravitational constraints are imposed Chandrasekaran:2022eqq ; Penington:2023dql ; AliAhmad:2023etg ; Klinger:2023tgi ; Kudler-Flam:2023qfl ; Jensen:2023yxy ; AliAhmad:2024eun ; Chen:2024rpx ; Kudler-Flam:2024psh ; Kaplan:2024xyk ; DeVuyst:2024grw ; Kolchmeyer:2024fly ; Penington:2024sum . This type reduction implies the emergence of a well-defined trace that is unique up to rescaling, enabling precise definitions of density matrices and von Neumann entropies, which are unavailable in the type III setting. The outcomes of this research program have been striking, including a new semiclassical definition of the generalized entropy Witten:2021unn ; Jensen:2023yxy ; Chandrasekaran:2022cip . In this type-reduction procedure, it is important to consider states admitting geometric and instantaneously-geometric modular flows, partially motivating our analysis.
@@
$S(\rho_{\Psi}^{\mathcal{V}})=-\alpha\int_{A}\text{d}\Sigma_{\mu}\,\frac{\xi^{\mu}_{\Psi}}{\norm{\xi_{\Psi}}^{d}}\,,$ (10)
where $\alpha$ is a state-independent thermal entropy density coefficient. This result is the covariant extension to higher dimensions of the $2d$ CFT expression in Wong:2013gua . We give various consistency checks of our formula in $d\geq 2$, including matching to several known examples. Notably, the physical reasoning which led us to propose eq. (10) is based on a local inverse temperature $\beta=2\pi\norm{\xi}$ associated instantaneously with observers along the flow, and thus we speculate it applies also to states with only instantaneously geometric modular flows.
Secondly, we define a procedure in $d=2$ for removing the entanglement between the left and right Rindler wedges in some state $|\Psi\rangle$. When $|\Psi\rangle=|\Omega\rangle$, this procedure yields the Rindler vacuum $|\Omega\rangle_{\mathcal{R}}$ [^7] This is the vacuum state with respect to one-sided boosts, rather than time-translations. It is disentangled, such that we can write $|\Omega_{\mathcal{R}}\rangle=|0\rangle_{\mathrm{left}}\otimes|0\rangle_{\mathrm{right}}$. By contrast, the usual Minkowski vacuum is entangled: $|\Omega\rangle=\sum_{n}e^{-\pi E_{n}}|n\rangle_{\mathrm{left}}\otimes|n\rangle_{\mathrm{right}}/\sqrt{Z}$, where $E_{n}$ is the eigenvalue of $|n\rangle_{\mathrm{right}}$ under right-sided boosts.. We then propose that in $d=2$, and for analogous disentangled states in higher $d$,
@@
In particular, this implies a general proposal for the Rindler vacuum energy-momentum tensor in CFTd. Again, the physical reasoning which leads to this formula is based on local temperatures associated instantaneously with observers along the flow.
@@
Various technical details are provided in the appendices. In appendix A, we demonstrate that if two states with geometric modular flows are related by a conformal unitary, the flows simply transform as the push-forward under the corresponding conformal map. In appendix B, we show that states with geometric modular flows in a subregion can be perturbed outside this subregion, without changing their flow inside, which is important for understanding the uniqueness of our various constructions. In appendix C, we ensure that the conformal mappings discussed in sections 2.1 and 2 are well-defined. In appendix D, we use the entanglement entropy formula discussed in Cardy:2016fqc to independently check the formula (61) in Wong:2013gua , which is a special case of our proposal eq. (10). Appendix E deals with nonlocal transformation rules for primary operators and the subtleties arising in the study of special conformal transformations in Minkowski spacetime. Finally, Appendix F shows that in the case of multiple intervals in CFT2, our proposed entropy formula (10) reproduces the correct entanglement entropy even when the modular Hamiltonian contains nonlocal terms. Readers seeking an introductory reference to modular flows and the theory of algebraic QFT are encouraged to consult Sorce:2023fdx ; Sorce:2024pte ; Sorce:2023gio ; Witten:2018zxz ; Hollands:2018wxj ; Hollands:2022dem ; Fewster:2019ixc ; Hollands:2014eia .
@@ 2 Excited states in $d=2$ with geometric modular flows
  • are future-directed in $\mathcal{V}$ and past-directed in $D(\bar{A})$, and timelike in both, [^9] When Haag duality holds, we have $U_{\Psi}^{D(A)}(s)=U_{\Psi}^{D(\bar{A})}(-s)$ Haag:1996hvx . Thus, if the modular flow with respect to $D(\bar{A})$ is future-directed in $D(\bar{A})$, then the modular flow with respect to $\mathcal{V}$ is past-directed in $D(\bar{A})$.
@@ 2.1 Excited states in the vacuum sector
The conformal Killing equation (8) in $d=2$ has general solution $\xi=c(u)\partial_{u}+b(v)\partial_{v}$. We accordingly define a vacuum-sector Unruh flow with respect to $\mathcal{V}=\mathcal{R}$ as an Unruh flow of this form, which approaches the vacuum modular flow Bisognano:1976za
@@
Let such a conformal transformation $\varphi$ be implemented at the quantum level by a conformal unitary $U_{\varphi}$. For a precise discussion of such unitaries, see e.g. Fewster:2004nj ; Oblak:2017ect . Under $U_{\varphi}$, conformal primary operators transform according to the first derivatives of $\tilde{u}(u)$, $\tilde{v}(v)$:
@@
Moreover, the Schwarzian appearing in the transformation of the stress tensor contains a third derivative, so will require that $\tilde{u}(u)$ and $\tilde{v}(v)$ are at least three times differentiable, i.e., $\tilde{u}(u),\tilde{v}(v)\in C^{3}(\mathbb{R})$. This guarantees that the state $|\Psi\rangle:=U_{\varphi}|\Omega\rangle$ has well-defined expectation values for most primary and quasi-primary operators of interest. [^13] Since higher derivatives will appear in transformation laws for descendants, infinite differentiability would be ideal, but it is difficult to explicitly check this in the cases of interest. Some of the conformal maps considered in this paper are not smooth, notably the disentangling maps of section 3.4; see further discussion in section 2. Smoothness, together with the falloff prescriptions (13), additionally helps ensure good behavior of the Bott-Thurston cocycle which defines the Virasoro central extension at the group level bott1977characteristic , e.g., see also Oblak:2017ect .
@@
$U_{\Psi}(s)=U_{\varphi}\,U_{\Omega}(s)\,U_{\varphi}^{\dagger}\,,$ (16)
because $U_{\Psi}(s)$ yields the relevant KMS condition. Note, eq. (16) holds also when the subregion is not preserved; see appendix A. The main point is that the modular flow for $|\Psi\rangle$ in $\mathcal{R}$ is a composition of geometric unitaries, and hence is geometric. [^14] Had we considered a highest-weight state $|h\rangle$ instead of the vacuum, the modular flow unitary $U_{\Psi}(s):=U_{\varphi}\,U_{h}(s)\,U_{\varphi}^{\dagger}$ (17) for $|\Psi\rangle:=U_{\varphi}|h\rangle$ has a geometric piece $U_{\varphi}$ and a generically non-geometric piece $U_{h}(s)$ (see, e.g., Caputa:2023vyr ).
@@
maps spatial infinity to the point $(\tilde{t}_{\cal{D}},\tilde{x}_{\cal{D}})=(0,R)$, such that $\mathcal{R}$ is mapped to $\mathcal{D}:=D(A)$, where $A$ is the interval $\tilde{x}_{\cal{D}}\in(-R,R)$ at $\tilde{t}_{\cal{D}}=0$. Since this map preserves the vacuum sector (see footnote 17), the pushforward of a vacuum-sector modular flow $\tilde{c}\,\partial_{\tilde{u}}+\tilde{b}\,\partial_{\tilde{v}}$ under this map is a vacuum-sector modular flow associated with the diamond $\mathcal{D}$. Since a generic state in the CFT is expected to have boost-like modular flow at $\partial A$, we expect
@@ 2.2 Excited states in the thermal sector
which exactly cancels eq. (37). This implies that there is the same amount of entropy in the two copies of the system, and thus the state is pure, as one can derive from eq. (69) below. As we derive later, in the thermal sector, the purity condition eq. (19) becomes
@@
These falloffs are more stringent than their analogue in eq. (13). [^19] Note that the arguments used in eq. (33) to fix the asymptotic behavior do not apply here, due to the absence of special conformal symmetry (see footnote 17). We excluded ${\cal O}(1)$ contributions, since the leading $\frac{\beta}{2\pi}$ term ensures the state reduces to the thermal state with inverse temperature $\beta$ at spatial infinity. We moreover excluded ${\cal O}(\frac{1}{\tilde{u}})$ contributions because otherwise, the resulting states would have infinitely more entanglement entropy in the Rindler wedge compared to $|\beta\rangle$, see section 3.2 for details. Note, however, that a similar analysis may apply for more general falloffs, as we discuss further in sections 3.2 and 6.
@@
What happens when an Unruh flow $\xi$ does not satisfy our purity condition? In that case, the vacuum modular flow can be transformed to $\xi$ using a conformal map $\varphi$ which fails to be differentiable at the origin. [^21] One can instead use a conformal map $\varphi$ which is not differentiable at infinity. Similar statements hold for flows not satisfying the boost-like property; then, $\varphi$ is non-invertible at the origin. See section 3.3. Due to its Schwarzian transformation rule (eq. (48) below), the resulting stress tensor is divergent along $\tilde{u}=0$ and $\tilde{v}=0$. These divergences propagate to other fields in the theory. [^22] Similarly, the Rindler vacuum stress tensor acquires divergences on the horizon due to the breaking of entanglement between left and right wedges Hartman:2015 ; Chen:2022 . The singular behavior resulting from violations of our purity condition reflects the idea that a pure state cannot be mapped to a mixed state by a unitary operator. [^23] For the vacuum sector, eq. (19) requires the entanglement entropy to be the same individually for both the left-moving and right-moving contributions. This indicates that our conformal unitaries cannot be used to introduce entanglement between right- and left-movers. Note that this is different from a shockwave state produced by the insertion of a primary operator $O_{1}$ at the origin; in this case, only primary operators with the same scaling dimension receive divergences at null separation.